OpenStudy (anonymous):

What is the answer of 1+1?

OpenStudy (bibby):

2

OpenStudy (anonymous):

Thank you sir.

OpenStudy (bibby):

Welcome to openstudy. You are VERY witty. Please read this: http://openstudy.com/code-of-conduct and enjoy your stay

OpenStudy (solomonzelman):

Please ask a serious question. This is a help site. DO you have anything to ask, that you are actually troubled with? If so, ask. And if not please close the question.

OpenStudy (solomonzelman):

\(\normalsize\color{blue}{ \frac{d}{dx}~~2x}\)

OpenStudy (solomonzelman):

this is the answer.

OpenStudy (anonymous):

solomon, pls you don't know calculus

OpenStudy (anonymous):

use this notation :D (2x)'

OpenStudy (solomonzelman):

No, I do not know calculus. The highest course I took is trig. I do know how to derive integrate and solve limits (in most of the cases)

OpenStudy (solomonzelman):

d/dx or ' . it is all same

OpenStudy (anonymous):

i dislike leibniz notation

OpenStudy (amistre64):

1+1=10

OpenStudy (solomonzelman):

Well, if you are deriving with respect to y, then you have to have a notation, no ?

OpenStudy (solomonzelman):

d/dx makes the equation look wiser sort of.

OpenStudy (anonymous):

yes in that case you would use leibniz, but with respect to x ' is better imo

OpenStudy (solomonzelman):

amistre, it is 11.

OpenStudy (bibby):

leibniz is traditional

OpenStudy (anonymous):

1+1 = window

OpenStudy (amistre64):

1+1+1 = 11

OpenStudy (solomonzelman):

that is 111

OpenStudy (amistre64):

thats 7 lol

OpenStudy (solomonzelman):

you put the 1s together.

OpenStudy (bibby):

leibniz is leibling leibniz is leben

OpenStudy (anonymous):

leibniz is lebanese